mathslover (mathslover):

Logarithm Question Find the number of zeroes between decimal point and first significant digit of \(\left(\cfrac{1}{3}\right)^{500}\) , if \(\log_{10} 3 = 0.47712\)

mathslover (mathslover):

This time, I promise, I don't have *any* idea of how to solve it.

mathslover (mathslover):

@Kainui @dan815 @ganeshie8 @BSwan

OpenStudy (dan815):

turn it into base 10 so u can see how many decimals its moving over

OpenStudy (anonymous):

hint :- \(log_b x= y\) \(b^y=x\)

OpenStudy (kainui):

\[\frac{1}{3}=3^{-1}\]\[3^{-500}\]\[\log_{10}3=.47712\]\[-500*\log_{10}3=-500*.47712\]\[3^{-500}=10^{-500*.47712}\]

mathslover (mathslover):

Okay, well, I almost got what @Kainui said. But, I'm really confused that how to get to the answer?

OpenStudy (dan815):

you got it alrready

OpenStudy (dan815):

oh kainui did

OpenStudy (kainui):

Well 10^(-1) means .1 and 10^(-2) means .01, so maybe that helps lead you into the right direction, since powers of 10 are literally decimal places now.

OpenStudy (dan815):

-500*0.477 = -238.5 so 238 decimals over s0 237 zeros

mathslover (mathslover):

10^{-238.56} yeah... got it....

mathslover (mathslover):

Thanks everyone! Specially, @Kainui , @BSwan and @dan815 .

OpenStudy (anonymous):

haha i did nothing :P

mathslover (mathslover):

Hmm, wouldn't it be 239 and 238 zeroes?

mathslover (mathslover):

238.56 \(\approx\) 239 \(10^{-239} \) ---> gives 238 zeroes between the decimal point and the first significant digit... right?

OpenStudy (dan815):

no

ganeshie8 (ganeshie8):

10^{-1} has 0 zeroes between decimal point and first sig fig 10^{-2} has 1 zero between decimal point and first sig fig 238-1 = 237

OpenStudy (dan815):

ok yes

OpenStudy (anonymous):

dnt use the sproimation 10^238.56= 10^238 . 10^ 0.56

OpenStudy (anonymous):

aproximation*

mathslover (mathslover):

@ganeshie8 - Yeah, the but what I think is : \(\boxed{\bf{500 \times 0.47712 = 238.56 \color{blue}{\approx 239}}}\) So, \(10^{-239} \) has 238 zeroes... That's what I think.

OpenStudy (dan815):

ya 239 decimal places and 239 zeros

OpenStudy (dan815):

238*

ganeshie8 (ganeshie8):

m feeling more dyslexic LOL yes 238 zeroes :)

OpenStudy (kainui):

http://www.wolframalpha.com/input/?i=10%5E-3.6 \[10^{-3.6} \approx .000251\] So instead of worrying about what the rule is or should possibly be at that really high number, let's just figure out the rule here. The rule is for our purposes, we round down from a decimal place apparently, that's the number of zeroes. He got the answer right a while ago, definitely 238.

OpenStudy (dan815):

go with what bswan said, because even if you approximated down it could still have 238 zeros

mathslover (mathslover):

Okay, yes, the book also says 238. Right!

OpenStudy (dan815):

like for example 10^-238.3 is still 238 zeros

mathslover (mathslover):

Oh, how ? \(10^{-238} \times 10^{-0.3} \) How will it have 238 zeroes?

OpenStudy (dan815):

yep because 10^ -0.3 is still 0.something right

OpenStudy (dan815):

as long as its anything more than 238

OpenStudy (dan815):

even 238.00000000000001

mathslover (mathslover):

Yeah, but, we needed to find the zeroes between the decimal point and the first significant digit.

mathslover (mathslover):

My calculator says : \(10^{-0.3} = 0.5011 ... \)

mathslover (mathslover):

So, if we consider the amount of zeroes : (betn the decimal point and first significant digit) we will come up with 237 , am I wrong with my concept here?

OpenStudy (dan815):

0.5*0.1=0.05

OpenStudy (dan815):

as long as number < 1 you will have 1 extra 0

mathslover (mathslover):

Oh! Got it... Thanks Dan.

OpenStudy (dan815):

ok

OpenStudy (dan815):

good :P